I am lost on this word problem. Would it be a polynomial to be factored or the product of two binomials? Thank you!
A 150-ft diagonal brace on a bridge connects a support of the center of the bridge to a side support on the bridge. The horizontal distance that it spans is 30 ft longer that the height that it reaches on the side of the bridge. Find the horizontal and vertical distances spanned by this brace.

Answer by mananth(16949)

The diagonal brace ,the horizontal & vertical support form a right angle
let height be x
horizontal distance = x+30
the brace is the diagonal
..
apply pythagoras theorem
x^2+(x+30)^2=150^2
x^2+x^2+60x+900=22500
2x^2+60x+900-22500=0
2x^2+60x-21600=0
/2
x^2+30x-10800=0
x^2+120x-90x-10800=0
x(x+120)-90(x+120)=0
(x+120)(x-90)=0
x=90 the positive value.
90 ft height & 120 ft horizontal span
...
m.ananth@hotmail.ca